OpenStudy (anonymous):

4y + x = 4 2y − x = 8 How do i get Y by itself?

OpenStudy (bloomlocke367):

well, you just have to follow out the order of operations to isolate it. For example, let's say you have 3x+6y=12. First, you'd have to subtract 3x from both sides to get 6y=12-3x. Then, you'd divide both sides by 6 to get y=-1/2x+2

OpenStudy (bloomlocke367):

Now, apply what I said to your problems.

OpenStudy (anonymous):

mine is simpler but yeah like s/he said

OpenStudy (anonymous):

so 4y - 4?

OpenStudy (anonymous):

yours made no sense @Thatman1

OpenStudy (bloomlocke367):

that would get x by itself, but didn't you ask for y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i need y by itself to chart it to find the intersect, i can do everything but get y by tself

OpenStudy (bloomlocke367):

okay. look at what I said and try again

OpenStudy (bloomlocke367):

brb.

OpenStudy (anonymous):

well i also need the 4 off is why im confused? its suppose to look like y = 4x + 4 (not the actual anwser i hope)

OpenStudy (bloomlocke367):

that is not the answer.

OpenStudy (anonymous):

well, yeah, but i just know how it looks before i graph it